### v2

# LoRE-Merging: Exploring Low-Rank Estimation For Large Language Model Merging

Abstract
While most current approaches rely on further training techniques, such as fine-tuning or reinforcement learning, to enhance model capacities, model merging stands out for its ability of improving models without requiring any additional training. In this paper, we propose a unified framework for model merging based on low-rank estimation of task vectors without the need for access to the base model, named LoRE-Merging. Our approach is motivated by the observation that task vectors from fine-tuned models frequently exhibit a limited number of dominant singular values, making low-rank estimations less prone to interference. We implement the method by formulating the merging problem as an optimization problem. Extensive empirical experiments demonstrate the effectiveness of our framework in mitigating interference and preserving task-specific information, thereby advancing the state-of-the-art performance in model merging techniques.
LoRE-Merging: Exploring Low-Rank Estimation For Large Language Model Merging
Zehua Liu1, Han Wu1, Yuxuan Yao2, Ruifeng She1, Xiongwei Han1 Tao Zhong1, Mingxuan Yuan1 1 Huawei Noah’s Ark Lab 2 Department of Computer Science, City University of Hong Kong liuzehua@connect.hku.hk wu.han1@huawei.com
1 Introduction
Large Language Models (LLMs) have become ubiquitous in numerous real-world applications (Bommasani et al., 2021; Zhuang et al., 2020). The utilization of LLMs typically involves fine-tuning them for specific tasks, a process that often yields superior performance compared to general-purpose LLMs. A rapidly emerging technique in this domain is model merging (Garipov et al., 2018; Wortsman et al., 2022; Yu et al., 2024b), which aims to create a single multi-task model by combining the weights of multiple task-specific models. This approach facilitates the construction of multi-task models by integrating knowledge from fine-tuned (FT) models without requiring additional training.
Building on recent studies (Ilharco et al., 2022; Yadav et al., 2024; Yu et al., 2024b), task vector-based merging approaches have demonstrated significant effectiveness, where task vectors are defined as the parameter differences between fine-tuned models and the base LLM. Achieving optimal results in model merging often requires minimizing interference between task vectors associated with different tasks. To address this, existing approaches utilize modified task vectors instead of the original ones. For instance, Yu et al. (2024b) applied random dropping with probability $p$ to obtain a sparse representation of task vectors, while Yadav et al. (2024) retained only the top-$k$ elements of each task vector based on magnitude, setting the remaining elements to zero. These strategies aim to produce sparse estimations of task vectors, a common technique for mitigating interference.
Nevertheless, task vector-based model merging approaches remain constrained by two fundamental limitations. First, the computation of task vectors necessitates access to the base model parameters and demonstrates heightened sensitivity to parametric variations (Yu et al., 2024b). As fine-tuning progress goes deeper, substantial parametric divergence emerges between the original base model and its fine-tuned counterpart, thereby greatly hindering them merging effectiveness (Yu et al., 2024a). Second, empirical evidence from Yadav et al. (2024) reveals that conflicting task vectors interactions could appear even when employing sparse estimation techniques. On the other hand, the sparsification process risks inadvertently eliminating essential task-specific features, thereby compromising the efficacy of the resultant merged model. These inherent constraints of sparse approximation methodologies underscore the necessity for developing alternative frameworks to estimate higher-fidelity low-rank task vector representations.
Figure 1: Singular value distributions for the task vector in layer $1$. We show the top-100 singular values, out of 4096 within the full rank.
To this end, we first empirically validate that task vectors exhibit a small number of dominant singular values, with the remaining singular values being significantly smaller in magnitude, as shown in Figure 1. Additionally, the dimension of the intersection of the images of two matrices is bounded by the minimum of their ranks. Therefore, we propose LoRE-Merging, a unified framework for model merging based on Low-Rank Estimation of task vectors, which eliminates the need for access to the base model. Specifically, given a set of FT models, we formulate the merging problem as an optimization problem whose goal is to simultaneously identify an approximate base model integrated with a set of low-rank task vectors. Together, these vectors collectively approximate the behavior of the FT models. By leveraging low-rank estimations, task vectors become inherently less susceptible to interference, effectively addressing a fundamental challenge in model merging. We conduct extensive experiments on optimization modeling problems and math word problems to confirm the effectiveness of our method.
2 Related Works
Merging fine-tuned models has been shown to offer several benefits, such as improving performance on a single target task (Gupta et al., 2020; Choshen et al., 2022; Wortsman et al., 2022), enhancing out-of-domain generalization (Cha et al., 2021; Arpit et al., 2022; Ilharco et al., 2022; Ramé et al., 2023), creating multi-task models from different tasks (Jin et al., 2022; Li et al., 2022; Yadav et al., 2024), supporting continual learning (Yadav and Bansal, 2022; Yadav et al., 2023), and addressing other challenges (Don-Yehiya et al., 2022; Li et al., 2022). Among these methods, task-vector-based merging approaches play an important role. Task Arithmetic (Ilharco et al., 2022) first introduced the concept of task vectors and shows that simple arithmetic operations can be performed to obtain the merged models. Building on this idea, methods like DARE (Yu et al., 2024b) and TIES-Merging (Yadav et al., 2024) adopt pruning-then-scaling techniques to merge task vectors, based on the assumption that not all parameters equally contribute to the final performance. However, these methods based on sparsity estimation consistently suffer from the interference among task vectors and require access to the base model, thus limiting their overall effectiveness.
3 Methodology
3.1 Problem Setting
We denotes $\mathcal{M}_{i}$ as the candidate models to be merged, where each $\mathcal{M}_{i}$ is parameterized by $\bm{\theta}_{i}$. In this work, we focus on the homogeneous model merging (Wortsman et al., 2022; Ilharco et al., 2022; Yadav et al., 2024), suggesting that the base models share the same model architecture. Specifically, these models can be obtained from the training process, such as checkpoints, or fine-tuned from the same pre-trained model, referred to as task-specific models. The primary objective of model merging is to construct a new model, $\mathcal{M}^{*}$, having better performance on the target single or multiple tasks.
3.2 Implicit Low-Rank Estimation for Model Merging
In this study, drawing upon methodologies similar to those presented by Matena and Raffel (2022), we investigate the model merging problem without presupposing specific characteristics of, or requiring access to, a base model. This methodological decision is underpinned by several key rationales. Firstly, in the context of checkpoint merging (Liu et al., 2024), a prevalent scenario involves access restricted solely to checkpoints saved during the training trajectory, before the finalization of a base model. Consequently, in such instances, the assumption of a pre-defined base model is untenable. Furthermore, as demonstrated by Yu et al. (2024b, a), model pairs frequently exhibit limited mergeability, particularly when subjected to extensive fine-tuning or prolonged pre-training, which can induce substantial parametric shifts. Under these circumstances, existing task-vector-based merging techniques often prove less effective due to significant representational divergence between an original base model and its fine-tuned counterpart. To surmount this challenge, we introduce LoRE-Merging, an implicit low-rank estimation approach to model merging. This method leverages the inherent robustness of low-rank estimation against perturbations while obviating the requirement for base model access.
The core idea of LoRE-Merging is straightforward: instead of using the original base model, we first construct an approximate base model and subsequently integrate the task-specific vectors via a low-rank approximation technique. Formally, denote the approximate base model as $\bm{\theta}_{0}$ and the low-rank task vectors $\{\bm{\delta}_{i}\}_{i=1}^{n}$ where $n$ is the number of FT models, our objective is to minimize the discrepancy between each FT model and its corresponding integrated version derived from the constructed base model, expressed as $\bm{\theta}_{0}+\bm{\delta}_{i}\approx\bm{\theta}_{i}$.
To ensure the low-rank structure of $\bm{\delta}$, we apply a nuclear norm penalty, as suggested in Cai et al. (2008). Then, we formulate the merging problem as the following optimization problem:
$\min_{\bm{\theta}_{0},\bm{\delta}_{1},\dots,\bm{\delta}_{n}}f:=\sum_{i=1}^{n}\left(\|\bm{\theta}_{0}+\bm{\delta}_{i}-\bm{\theta}_{i}\|_{F}^{2}+\mu\|\bm{\delta}_{i}\|_{*}^{2}\right),$ (1)
where $\|\cdot\|_{*}$ represents the nuclear norm, and $\mu>0$ is a hyperparameter. In Equation (1), the first term minimizes the difference between $\bm{\theta}_{0}+\bm{\delta}_{i}$ and $\bm{\theta}_{i}$, ensuring reconstruction accuracy. The second term acts as a penalty that encourages the task vectors $\bm{\delta}_{i}$ to exhibit low-rank properties.
This problem is a standard multi-variable convex optimization problem. To solve it efficiently, we employ the coordinate descent method (Wright, 2015). Starting from an initial point $\{\bm{\theta}_{0}^{0},\bm{\delta}_{1}^{0},\dots,\bm{\delta}_{n}^{0}\}$, each iteration (round $k+1$) updates the variables by iteratively solving the following single-variable minimization problem:
$\begin{dcases}\bm{\theta}_{0}^{k+1}=\mathop{\arg\min}_{\bm{\theta}}f(\bm{\theta},\bm{\delta}_{1}^{k},\cdots,\bm{\delta}_{n}^{k})\\ \bm{\delta}_{i}^{k+1}=\mathop{\arg\min}_{\bm{\delta}}f(\cdots,\bm{\delta}_{i-1}^{k},\bm{\delta},\bm{\delta}_{i+1}^{k},\cdots),~{}\forall i\end{dcases}$ (2)
The update for $\bm{\theta}_{0}^{*}$ is trivial, while the update for $\bm{\delta}$ is less straightforward due to the presence of the nuclear norm. Fortunately, as shown in Cai et al. (2010), closed-form solutions for the coordinate descent method iterations in Problem (1) can be obtained using the Singular Value Thresholding (SVT) technique. Recall that for a given matrix $\bm{\delta}$ with the Singular Value Decomposition (SVD) $\bm{\delta}=U\Sigma V^{\top}$, and a hyperparameter $\mu$, the SVT operator is defined as follows. Let $\Sigma^{+}(\mu):=\text{diag}((\sigma_{i}-\mu)^{+})$, where $(\cdot)^{+}$ denotes the positive part function. The SVT($\bm{\delta};\mu$) operator with hyperparameter $\mu$ is then defined as SVT($\bm{\delta};\mu$) := $U\Sigma^{+}(\mu)V^{\top}$. Using the SVT operator, the update for $\bm{\delta}_{i}$ can be expressed as: $\bm{\delta}_{i}^{k+1}=$ SVT($\bm{\theta}_{i}-\bm{\theta}_{0}^{k+1};\mu$).
Once the optimization problem is solved, we can obtain the approximate base model and a set of low-rank task vectors. Then, existing task-vector based approaches, such as Average Merging and TIES-Merging, can be applied to combine the task vectors and the base model. In this work, we directly adopt Average Merging as our post-calculation merging methods for simplicity, as as it demonstrated comparable performance to TIES-Merging in our preliminary experiments. The overall process is outlined in Algorithm 1.
4 Experiments
Method DPSK & Numina LM & Math Math & Code Checkpoints Merging Avg.
GSM8K MATH GSM8K MATH MMLU GLUE MBPP EasyLP ComplexLP NL4OPT
Baseline 76.3 55.8 54.8 12.4 52.0 63.3 32.0 81.9 39.3 94.0 56.18
Average 75.0 45.8 58.8 12.6 52.8 61.7 28.0 75.9 40.3 91.6 54.25
DARE 81.0 54.2 14.9 3.7 52.7 59.1 27.6 80.7 35.1 95.1 50.41
TIES 80.8 51.6 58.5 11.8 53.1 59.3 26.8 82.4 42.7 94.8 56.18
[…]
Table 1: Evaluations on various benchmarks. LM and Math are Wizard-series models, namely WizardLM-13B and WizardMath-13B. Code is llama-2-13b-code-alpaca model. The score of baseline is the higher one of base models.
Datasets $\mu=0$ $\mu=0.01$ $\mu=0.1$ $\mu=1.0$
GSM8K (%) 81.3 82.0 79.9 67.3
MATH (%) 53.8 54.5 53.8 42.4
Table 2: The ablation study for the hyperparameter $\mu$ (with $\lambda=1.0$) on DPSK & Numina.
Datasets $\lambda=0.5$ $\lambda=1.0$ $\lambda=1.5$
GSM8K (%) 18.9 82.0 79.1
MATH (%) 33.1 54.5 51.0
Table 3: The ablation study for the hyperparameter $\lambda$ (with $\mu=0.01$) on DPSK & Numina.
Baselines & Settings
We compare LoRE-Merging with following popular merging methods. Average Merging (Choshen et al., 2022): This method computes the element-wise mean of all the individual models. DARE (Yu et al., 2024b): This approach randomly drops task-specific vectors and rescales the remaining vectors back to the base model. We set the hyperparameter for the random probability to $0.5$. TIES-Merging (Yadav et al., 2024): In this method, task-specific vectors are randomly dropped, and only the parameters aligned with the final agreed-upon sign are merged. For TIES-merging, we set the top-$k$ value to $20\%$, and the hyperparameter $\lambda$ is fixed at $1$. For LoRE-Merging, the rank $r$ is determined dynamically. For a given task vector $\bm{\delta}\in\mathbf{R}^{m\times n}$, we set the rank $r=0.2\times\min\{m,n\}$ to get a low-rank estimation.
Evaluation
We first evaluate LoRE-Merging on diverse benchmarks, including GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., ) (math word problem), MMLU (Hendrycks et al., ), GLUE(Wang et al., 2019) (commonsense reasoning) and MBPP(Austin et al., 2021) (code task). We evaluate DeepSeek-series models (NuminaMath-7B (Beeching et al., 2024) and DeepSeek-Math-7B-Base (Shao et al., 2024)) and LLaMA-series models (WizardLM-13B (Xu et al., 2023), WizardMath-13B (Luo et al., 2023) and LLaMA-2-13B-Code model). Additionally, we also evaluate on the advanced task, i.e. mathematical optimization modeling problems (Ramamonjison et al., 2023; Huang et al., 2024, 2025). This task aims to generate solvable mathematical models given an optimization problem in natural language. As the lack of public models on this task, we first fine-tuned Qwen-2.5-Coder-7B-Instruct model (Hui et al., 2024) with the dataset provided by Huang et al. (2025) and merge checkpoints in the training process. The evaluations are conducted on MAMO dataset (Huang et al., 2024) which includes two subsets EasyLP and ComplexLP, and NL4OPT dataset (Ramamonjison et al., 2023).
Main Results
As shown in Table 1, LoRE-Merging achieves superior performance across most metrics, as well as the highest overall score. For the math word problem evaluation, our method demonstrates consistently superior performance against baselines, except for the evaluations on MATH (DPSK & Numina) and MBPP datasets. We think this is because of the significant performance gap between the base models, where DeepSeek-Math achieves only a score of 36.2 on the MATH dataset, while NuminaMath reaches 55.8. As indicated in Yao et al. (2024), a large performance gap can significantly impact the effectiveness of model merging. Another worthy-noting observation is that DARE demonstrates significantly poorer performance when merging WizardLM and WizardMath. This can likely be attributed to the substantial parameter divergence between these models, which results in the failure of calculating the task vector derived from the base model. In contrast, our LoRE-Merging with the approximate base model and low-rank task vectors demonstrates superior robustness and effectiveness in solving math word problems. For the evaluations on optimization modeling with checkpoints merging, we can see existing task-vector based merging methods consistently improve the performance because of the marginal gap between the checkpoints. Therefore, we believe that checkpoint merging can serve as a highly effective technique complementary to training methods, particularly our LoRE-Merging method. We also conduct a detailed analysis how our method enhance the modeling capacity on ComplexLP dataset. We found that the earlier checkpoint is more good at identifying the variables and parameters in the questions while the later one focuses on more complex components, such as formulating variables and the constraints. With the merging of task vectors, the merged model exhibits superior overall performance on the task.
Ablations
We conduct a systematic empirical analysis of the selection of hyperparameters $\lambda$ and $\mu$, as presented in Table 3 and Table 3. Our results show that the best performance is achieved with $\lambda=1.0$ and $\mu=0.01$. Notably, variations in the hyperparameters around these values do not significantly impact the final performance, indicating the robustness of LoRE-Merging.
5 Conclusion
In this paper, we propose a unified framework for merging models based on low-rank estimation, named LoRE-Merging. We achieve it by formulating the merging problem as an optimization problem. Extensive experiments demonstrate the efficacy and efficiency of our proposed methods.
Limitations
Although we have demonstrated the effectiveness of our method on merging homogeneous models, we have not yet evaluated it on merging heterogeneous models which is a much more challenging task. Compared to existing task-vector based model merging methods, our method is the most suitable one that can be adapted to heterogeneous model merging, as we disentangle the base model and task vectors. We think how to expand LoRE-Merging to heterogeneous model merging should be a promising future direction.
Appendix A Appendix
Method Average TIES-Merging Twin-Merging LoRE-Merging
Acc. on GSM8K 75.0 80.8 79.9 81.0
Runtime 4.2s 5min 29s 17min 44s 12min 24s
Table 4: Time cost comparisons among different merging methods.
A.1 Speed and Computational Cost
While standard SVD exhibits computational inefficiency for extremely large matrices comprising billions of elements, its application to LLM presents a substantially different computational profile. Despite LLMs containing billions of parameters in aggregate, SVD operations are performed on individual parameter matrices, each typically comprising only millions of entries. For instance, in the Qwen2.5-72B architecture, the largest matrix requiring decomposition is dimensioned at $8192\times 28564$, while for Qwen2.5-7B, the corresponding matrix has dimensions of $3854\times 18944$. Thus, the substantial parameter differential between LLM scales does not translate to proportionally expanded matrix dimensions. In our implementation, merging operations for 7B-scale models require approximately 5 minutes using Ties-Merging, while LoRE-Merging necessitates approximately 12 minutes. However, compared to another SVD-based mering method, like Twin-Merging (Lu et al., 2024), our method exhibit superior performance on efficiency.
A.2 Task Vector Rank Validation
In this subsection, we validate the low-rank properties underlying the low-rank assumption. Specifically, we focus on the checkpoint merging problem and compute the rank of the task vectors. As previously discussed, we set the rank $r$ as $r=0.2\times\min\{m,n\}$ for any given task vector $\bm{\delta}$.
The distribution of the largest 100 singular values for Layer $1$ is presented in Figure 1. Our experimental results reveal that $\sigma_{r}\leq 0.05\times\sigma_{1}$, indicating that the singular values set to 00 in low-rank estimation are significantly smaller than the largest singular value across all linear layers. This finding supports the validity of adopting a low-rank approximation for task vectors, as it reflects the inherent structure of the data.
Algorithm 1 Implicit low-rank merging method
Input: fine-tuned models $\{\bm{\theta}_{i}\}_{i=1}^{n}$, parameter dimension $d$, and hyperparameter $\lambda,\mu$.
Output: merged model $\bm{\theta}^{*}$.
[…]
  Set $\bm{\delta}_{i}=0$ for $i=1,2,\dots,n$.
  while iteration NOT converges do
     $\bm{\theta}_{0}=\frac{1}{n}\sum_{i=1}^{n}(\bm{\theta}_{i}-\bm{\delta}_{i})$
     for $i=1,\dots,n$ do
        $\bm{\delta}_{i}=\mathrm{SVT}(\bm{\theta}_{i}-\bm{\theta}_{0};\mu)$;
     end for
  end while
[…]
  $\bm{\tau}=\sum_{i=1}^{n}\bm{\delta}_{i}$.
[…]
  $\bm{\gamma}=sgn(\sum_{i=1}^{n}\bm{\delta}_{i})$.
  for $p=1,2,\dots,d$ do
     $\mathcal{A}^{p}=\{i:\bm{\gamma}_{i}^{p}=\bm{\gamma}^{p}\}$
     $\bm{\tau}^{p}=\frac{1}{|\mathcal{A}^{p}|}\sum_{i\in\mathcal{A}^{p}}\bm{\tau}^{p}$
  end for
[…]
  $\bm{\theta}^{*}=\bm{\theta}_{0}+\lambda\bm{\tau}$.
  return  $\bm{\theta}^{*}$